OpenStudy (anonymous):

Which operation is equivalent to dividing by 1/5?

OpenStudy (anonymous):

multiplying times 5.

OpenStudy (luigi0210):

Multiplying by 5.. \[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}*\frac{d}{c} \]

OpenStudy (anonymous):

\[\div \frac{A}{B}~~~->~~~\times \frac{B}{A}\]